# code[Code] DistMLIP: A Distributed Inference Platform for Machine Learning Interatomic Potentials

### v1

@@ Abstract
Large-scale atomistic simulations are essential to bridge computational materials and chemistry to realistic materials and drug discovery applications. In the past few years, rapid developments of machine learning interatomic potentials (MLIPs) have offered a solution to scale up quantum mechanical calculations. Parallelizing these interatomic potentials across multiple devices poses a challenging, but promising approach to further extending simulation scales to real-world applications. In this work, we present DistMLIP, an efficient distributed inference platform for MLIPs based on zero-redundancy, graph-level parallelization. In contrast to conventional space-partitioning parallelization, DistMLIP enables efficient MLIP parallelization through graph partitioning, allowing multi-device inference on flexible MLIP model architectures like multi-layer graph neural networks. DistMLIP presents an easy-to-use, flexible, plug-in interface that enables distributed inference of pre-existing MLIPs. We demonstrate DistMLIP on four widely used and state-of-the-art MLIPs: CHGNet, MACE, TensorNet, and eSEN. We show that existing foundational potentials can perform near-million-atom calculations at the scale of a few seconds on 8 GPUs with DistMLIP.
@@ 1 Introduction
Atomistic simulation has been the workhorse in computational materials and drug discovery over the recent years (Merchant et al., 2023; Jain et al., 2013; De Vivo et al., 2016). The chemical properties and behavior of a material are essentially determined by the interactions in the given set of atomic arrangements. In the most simplified framework, one can formulate this problem as solving the function that determines the potential energy surface (PES) of a set of atoms given by , where is the energy, and are the positions and chemical identities of the atoms.
@@
To study the material’s properties, multiple fundamentally different methods have been developed to obtain or construct the function . Empirical interaction rules such as the electrostatic interactions and Lennard-Jones potential (Jones, 1924) qualitatively describe the bond energy between atoms, but only predict the very fundamental behavior of the materials. Classical force fields (FF) like embedded atom methods (Daw and Baskes, 1984), CHARM (Vanommeslaeghe et al., 2010), and Amber (Wang et al., 2004) improved the prediction of the PES by adding more descriptor terms into . These classical FFs are cheap, intuitive, and explainable, but are often not accurate enough and have been constructed and applied to narrow chemical domains and a few elements. Fundamentally, the interactions in are determined by the electronic structure of a material and can be solved in first principles by quantum mechanics. Quantum chemical simulation methods, such as Density Functional Theory (DFT) (Perdew et al., 1996) and coupled clustering (CC) methods (Raghavachari et al., 1989), enabled the ab initio calculations of atomic behavior that are much more accurate than empirical methods. However, their computational complexity limits the practical use of quantum chemical simulation methods for many realistic applications. DFT, the most widely used ab initio simulation method, scales cubically with the number of electrons and is therefore limited to simulating only a few hundred atoms (Beck, 2000). Prohibitively high computational cost makes DFT only useful in describing materials properties that can be learned from a small simulation cell (Wang et al., 2024).
@@
Machine learning approaches such as machine learning interatomic potentials (MLIPs) open the possibility to increase simulation scale while retaining quantum chemical accuracy by building ML surrogate models trained on DFT and CC data (Wang et al., 2024; Gasteiger et al., 2021; Deng et al., 2023; Batzner et al., 2022; Musaelian et al., 2023; Fu et al., 2025). Compared to feature-based classical FFs, deep-learning-based MLIPs enable improved learnability to better model the PES data. Graph neural networks (GNNs), especially, have demonstrated extraordinary computational efficiency and accuracy by learning both long-range and high-order atomic interactions through message passing. By design, the computation time of MLIPs scales linearly with the number of atoms , enabling simulations with tens of thousands of atoms at nano-second time scale.
@@
Many materials engineering problems involve finite-sized effects like protein folding (Jumper et al., 2021), interfacial reactions (Du et al., 2023), particle-size effects (Shi et al., 2020), and formation of nano-domains (Holstun et al., 2025). Such systems require meso-scale simulations with upwards of millions of atoms, necessitating the ability to further scale the capacity of MLIP simulations. One promising solution is to expand MLIP inference from single-device to multi-device inference. Simulation packages such as LAMMPS provide ad hoc solutions for multi-GPU simulation (Thompson et al., 2022). Multi-device simulations are realized by dividing the total simulation cell into multiple, mutually exclusive, small cells for each device. Each small cell is then padded with additional atoms beyond the cell boundary in order to properly calculate the energy and forces within the small cell. This method, known as spatial partitioning, is based on the fundamental assumption that the force field only contains short-range interactions (Plimpton, 1995).
@@
Since most MLIPs are designed to be long-ranged, expanding the number of utilized GPUs during inference time is a nontrivial task due to the necessity to distribute the large system cell across multiple devices. Currently, there exists no native multi-GPU support for GNN-based MLIPs as most MLIPs have been implemented for only single-device inference. In order to support fast, accurate, and parallelizable atomistic simulations, we hereby present a distributed MLIP inference platform, DistMLIP, that enables efficient multi-device inference without the need for a modified architecture or additional training. Our highlighted contributions are as follows:
@@
DistMLIP features an efficient, general, and versatile parallel inference platform for MLIP inference. By design, most popular MLIPs can be supported with a minimal amount of adaptation. In this work, we include benchmarking results of 4 widely used MLIPs: MACE, TensorNet, CHGNet, and eSEN.
@@
DistMLIP supports the calculation of stress in distributed inference, which allows isobaric molecular dynamics (MD) simulations with varying cell sizes.
To allow flexible usage, our implementation does not depend on a 3rd party distributed simulation library such as LAMMPS. As a result, DistMLIP supports plug-in usage of any MLIP workflow.
All code in DistMLIP is open sourced at https://github.com/AegisIK/DistMLIP.
@@ 2.1 Machine Learning Interatomic Potentials
Machine learning interatomic potentials (MLIPs) predict the potential energy surface (PES) by inputting atomic numbers, atomic positions, and optionally the lattice, and outputting the energy and forces of the system (Bartók et al., 2010; Zhang et al., 2018). The most common MLIP architecture today is the graph neural network (GNN), where nodes in the atom graph represent atoms and edges in the atom graph represent the pair-wise distances between atoms that are within a pre-defined cutoff distance (Batzner et al., 2022; Simeon and De Fabritiis, 2023; Passaro and Zitnick, 2023; Gasteiger et al., 2021; Schütt et al., 2021, 2018; Smith et al., 2017). GNN computation scales linearly with the number of atoms, as the amount of computation is associated with the neighbors within the receptive field of each atom. Some MLIPs also pass messages on top of higher-order graphs, such as threebody bond graphs, that encode angles as pairwise information between bonds (Choudhary and DeCost, 2021; Deng et al., 2023; Yang et al., 2024). MLIPs built on top of the transformer architecture have also been introduced, where "tokens" represent individual nodes and full self attention is performed over all tokens (Liao et al., 2024; Vaswani et al., 2017). Recently, a class of foundational potentials have been shown to generalize across diverse chemistries by pretraining on massive datasets (Chen and Ong, 2022; Deng et al., 2023; Chanussot* et al., 2021; Barroso-Luque et al., 2024; Yang et al., 2024; Merchant et al., 2023; Kaplan et al., 2025). These pretrained foundational potentials substantially reduce the need for target-system training, and their open-sourced pretrained checkpoints serve as ready-to-use universal MLIPs.
@@ 2.2 Spatial Partitioning
LAMMPS implements multi-GPU inference via a spatial partitioning approach where the simulation space is split into mutually exclusive partitions. For each mutually exclusive partition, LAMMPS creates a second, larger partition that includes all atoms up to the interaction radius of the FF, commonly known as border or ghost nodes. This is required as the energy and force calculation of the atoms within each mutually exclusive partition requires the atomic information from all nodes within the model’s interaction radius. This leads to highly redundant calculations as the computation performed on the ghost nodes is thrown away after each time step. By estimate, a 64-molecule water system calculated with a 6-layer GNN that has a 6 angstrom cutoff distance would require the computation of 20,834 ghost atoms when using spatial partitioning (Musaelian et al., 2023). Furthermore, unlike classical FFs, most current MLIPs do not have a mature interface with LAMMPS, making spatial partitioning practically infeasible for the majority of MLIPs that have been developed.
@@
DeepMD has been applied to the simulation of 100 million atoms of water by 27,360 NVIDIA V100 GPUs on the Summit supercomputer utilizing the spatial partitioning features within LAMMPS Jia et al. (2020). The atomic system size was further extended to 10 billion atoms after further optimization of model tabulation, kernel fusion, and redundancy removal of the Deep Potential architecture (Guo et al., 2022).
@@
Allegro has been developed as a strictly local, E(3)-equivariant interatomic potential that features efficient parallelization through space partitioning due to its short-range design (Musaelian et al., 2023). Instead of performing message passing where messages are passed between nodes several times, it encodes each atom-pair’s relative geometry in learned spherical harmonic components and captures many-body interactions via trainable tensor products of these features. Because of this strict locality, Allegro scales particularly well on large atomic systems – Kozinsky et al. (2023) used Allegro and LAMMPS to simulate a bulk Ag model with 100 million atoms, achieving 0.003 microseconds/atom-timestep using 128 NVIDIA-A100-80GBs.
@@
However, strictly local models experience key limitations. The need for efficient parallelization restricts their interaction range to only a few angstroms, preventing their use on systems that require the modeling of long-range interactions (Zhou et al., 2023; Song et al., 2024; Gong et al., 2025; Cheng, 2025; Anstine and Isayev, 2023). Furthermore, the short-range design prevents the MLIP’s application from simultaneously learning diverse chemical environments. As the short-range MLIP’s cutoff is often determined by the radial distribution function of one targeted material system, it is infeasible to determine a universal cutoff that works for many materials if the training dataset contains more than one chemical system, which is becoming a common scenario with the increased interest in foundational interatomic potentials. These problems raise the need for a simple, unified, and versatile API to parallelize MLIPs.
2.3 Graph Partitioning
Graph partitioning algorithms find applications in solving PDEs via domain decomposition, solving sparse linear systems of equations, circuit partitioning and layout, VLSI design, social network analysis, clustering algorithms, and image segmentation (Pothen, 1997; Stanton and Kliot, 2012; Bader et al., 2013; Tolliver and Miller, 2006; Peng et al., 2013). One common use case in graph partitioning is to create mutually exclusive spanning sets of nodes that contain the minimum number of edges that cross from one partition to another. This use case can be tackled using sparse, symmetric matrix reordering methods such as the Reverse Cuthill-Mckee (RCM) algorithm, which permutes a sparse matrix to minimize its bandwidth (i.e. reordering rows and columns such that non-zero values are closer to the diagonal) (Cuthill and McKee, 1969; Azad et al., 2017). The reordered matrix can then be partitioned along the columns in order to calculate graph node partitions. METIS is a graph partitioning algorithm that utilizes a graph coarsening phase, an initial partitioning sequence over the coarsened graph, and an uncoarsening and partition refinement stage (Karypis and Kumar, 1998). However, applications depending on algorithms such as RCM or METIS typically don’t have latency requirements during the graph partitioning stage. In atomistic simulation, the underlying graphs are recalculated and repartitioned at each time step. Therefore, even small latency increases during the graph creation and partitioning stage get compounded into significant increases in overall simulation time. In our own experiments, we find that RCM and METIS could increase inference time for million-atom graphs by several seconds per timestep.
@@ 3 Methods
Figure 1(a) denotes the usage of DistMLIP. Public MLIP models can be easily adapted to perform distributed large simulations with DistMLIP. The core infrastructure of DistMLIP is in the construction of graphs, subgraphs, and communication-related metadata.
@@ 3.1 Graph-Parallel Message Passing
After the material system is converted into a graph using a neighbor list construction algorithm, it can be partitioned into subgraphs for each device, as illustrated in Fig. 1(b). At each graph convolution, each node’s features are updated according to the edge and node features of its incoming neighbors. We can partition the nodes of the graph into disjoint sets, constructing graphs , where is the number of partitions. Each of the graph partitions are distributed to its own GPU. To accurately calculate the features after one graph convolution, we expand into , where consists of all nodes such that there exists an edge with . Formally,
@@
After each graph convolution, we transfer the border node and border edge features to and from each partition, as shown in Fig. 1(d). After this transfer process, each GPU has the most updated node and edge feature to begin the next convolution. This implementation is completely model-agnostic and can be applied to both conservative and direct force prediction interatomic potentials.
@@ 3.2 Distributing Atom Graphs
3.3 Distributing Three-body Graphs
Distributing the bond graph involves selecting all 1-hop and 2-hop neighbors of the pure atom graph nodes assigned to a partition. We then create an edge table mapping from node ids to edges originating from the node id pointing to a different node. By recursively traversing the table, we are able to create the bond graph for each partition in parallel. Border nodes within the bond graph are associated with the 1-hop edge neighbors of border edges within the atom graph – hence necessitating the inclusion of 2-hop neighbors. The parallel bond graphs thus contain the 1-hop neighbors of each pure bond graph node assigned to the partition. The complete procedure is found in algorithm 2 of the appendix.
@@ 3.4 Currently Supported MLIPs
Currently, we have implemented four widely used models in DistMLIP: 1) CHGNet (Deng et al., 2023), an invariant MLIP that features both an atom graph as well as a bond graph, 2) TensorNet (Simeon and De Fabritiis, 2023; Ko et al., 2025), an atom graph-only, computationally-efficient, equivariant MLIP with performance on par with Nequip, 3) MACE (Batatia et al., 2023), an atom graph-only equivariant MLIP that directly models many body interactions between atoms, and 4) eSEN (Fu et al., 2025), an atom graph-only, smooth and equivariant MLIP. Specific architecture details are found in appendix C.
3.5 Usage

### v2

@@ Abstract
Large-scale atomistic simulations are essential to bridge computational materials and chemistry to realistic materials and drug discovery applications. In the past few years, rapid developments of machine learning interatomic potentials (MLIPs) have offered a solution to scale up quantum mechanical calculations. Parallelizing these interatomic potentials across multiple devices poses a challenging, but promising approach to further extending simulation scales to real-world applications. In this work, we present DistMLIP, an efficient distributed inference platform for MLIPs based on zero-redundancy, graph-level parallelization. In contrast to conventional spatial partitioning parallelization, DistMLIP enables efficient MLIP parallelization through graph partitioning, allowing multi-device inference on flexible MLIP model architectures like multi-layer graph neural networks. DistMLIP presents an easy-to-use, flexible, plug-in interface that enables distributed inference of pre-existing MLIPs. We demonstrate DistMLIP on four widely used and state-of-the-art MLIPs: CHGNet, MACE, TensorNet, and eSEN. We show that DistMLIP can simulate atomic systems 3.4x larger and up to 8x faster compared to previous multi-GPU methods. We show that existing foundation potentials can perform near-million-atom calculations at the scale of a few seconds on 8 GPUs with DistMLIP.
@@ 1 Introduction
Atomistic simulation has been the workhorse in computational materials and drug discovery over the recent years (Merchant_2023; Jain_2013; role_drugs). The chemical properties and behavior of a material are essentially determined by the interactions in the given set of atomic arrangements. In the most simplified framework, one can formulate this problem as solving the function that determines the potential energy surface (PES) of a set of atoms given by , where is the energy, and are the positions and chemical identities of the atoms.
@@
To study the material’s properties, multiple fundamentally different methods have been developed to obtain or construct the function . Classical force fields (FF) like embedded atom methods (EAM1984), CHARMM (CHARMM_2010), and Amber (Amber_2004) qualitatively predict the PES and the bond energy between atoms. These classical FFs are cheap, intuitive, and explainable, but are often not accurate enough and have been constructed and applied to narrow chemical domains and a few elements. Fundamentally, the interactions in are determined by the electronic structure of a material and can be solved in first principles by quantum mechanics. Quantum chemical simulation methods, such as Density Functional Theory (DFT) (Perdew_Burke_Ernzerhof_1996) and coupled cluster (CC) methods (Raghavachari_Trucks_Pople_Head-Gordon_1989), enabled the ab initio calculations of atomic behavior that are much more accurate than empirical methods. However, their computational complexity limits the practical use of quantum chemical simulation methods for many realistic applications. DFT, the most widely used ab initio simulation method, scales cubically with the number of electrons and is therefore limited to simulating only a few hundred atoms (beck2000real). Prohibitively high computational cost makes DFT only useful in describing materials properties that can be learned from a small simulation cell (ai2bmd).
@@
Machine learning approaches such as machine learning interatomic potentials (MLIPs) open the possibility to increase simulation scale while retaining quantum chemical accuracy by building ML surrogate models trained on DFT and CC data (Bartók2010GAP; zhang2018deeppotential; ai2bmd; gemnet; deng2023chgnet; Batzner_2022_Nequip; Musaelian_2023; esen). Compared to feature-based classical FFs, deep-learning-based MLIPs enable improved learnability to better model the PES data. Graph neural networks (GNNs), especially, have demonstrated extraordinary computational efficiency and accuracy by learning both long-range and high-order atomic interactions through message passing. By design, the computation time of MLIPs scales linearly with the number of atoms , enabling simulations with tens of thousands of atoms at nano-second time scale.
@@
Many materials engineering problems involve finite-sized effects like protein folding (Jumper2021AlphaFold), interfacial reactions (Du2023SurfaceReconstruct), particle-size effects (Shi2020ParticleSize), and formation of nano-domains (Holstun2025NanoDomain). Such systems require meso-scale simulations with upwards of millions of atoms, necessitating the ability to further scale the capacity of MLIP simulations. One promising solution is to expand MLIP inference from single-device to multi-device inference. Simulation packages such as LAMMPS provide ad hoc solutions for multi-GPU simulation (lammps). Multi-device simulations are realized by dividing the total simulation cell into multiple, mutually exclusive, small cells for each device. Each small cell is then padded with additional atoms beyond the cell boundary in order to properly calculate the energy and forces within the small cell. This method, known as spatial partitioning, is based on the fundamental assumption that the force field only contains short-range interactions (Plimpton1995ParralelMD).
@@
Since most MLIPs are designed to be relatively long-ranged, expanding the number of utilized GPUs during inference time is a nontrivial task due to the necessity to distribute the large system cell across multiple devices. Currently, there exists no native multi-GPU support for GNN-based MLIPs as most MLIPs have been implemented for only single-device inference. In order to support fast, accurate, and parallelizable atomistic simulations, we hereby present a distributed MLIP inference platform, DistMLIP, that enables efficient multi-device inference without the need for a modified architecture or additional training. Our highlighted contributions are as follows:
@@
DistMLIP features a simple, efficient, general, and versatile parallel inference platform for MLIP inference. By design, most popular MLIPs can be supported with a minimal amount of adaptation. In this work, we include benchmarking results of 4 widely used MLIPs: MACE, TensorNet, CHGNet, and eSEN.
@@
To allow flexible usage, DistMLIP does not depend on a 3rd party distributed simulation library such as LAMMPS. As a result, DistMLIP supports plug-in usage of any MLIP workflow.
We show that the simple-yet-effective partitioning technique DistMLIP utilizes performs MD up to 8x faster compared with the more standard graph partitioning techniques.
@@ 2.1 Machine Learning Interatomic Potentials
The most common MLIP architecture today is GNN, where nodes in the atom graph represent atoms and edges in the atom graph represent the pair-wise distances between atoms that are within a pre-defined cutoff distance (Batzner_2022_Nequip; tensornet; escn; gemnet; painn; schutt2018schnet; smith2017ani). GNN computation scales linearly with the number of atoms, as the amount of computation is associated with the neighbors within the receptive field of each atom. Some MLIPs also pass messages on top of higher-order graphs, such as threebody bond graphs, that encode angles as pairwise information between bonds (Choudhary_alignn_2021; deng2023chgnet; yang2024mattersim). MLIPs built on top of the transformer architecture have also been introduced, where "tokens" represent individual nodes and full self-attention is performed over all tokens (liao2023equiformerv2; vaswani2017attention). Recently, a class of foundation potentials (FPs) have been shown to generalize across diverse chemistries by pretraining on massive datasets (Chen_Ong_2022; deng2023chgnet; oc20; omat; yang2024mattersim; Merchant_2023; Kaplan2025MatPES). These pretrained FPs substantially reduce the need for target-system training, and their open-sourced pretrained checkpoints serve as ready-to-use universal MLIPs.
@@ 2.2 Spatial Partitioning
LAMMPS implements multi-GPU inference via a spatial partitioning approach where the simulation space is split into mutually exclusive partitions. For each mutually exclusive partition, LAMMPS creates a second, larger partition that includes all atoms up to the interaction radius of the FF, commonly known as border or ghost nodes. This is required as the energy and force calculation of the atoms within each mutually exclusive partition requires the atomic information from all nodes within the model’s interaction radius. This leads to highly redundant calculations as the computation performed on the ghost nodes is thrown away after each time step. By estimate, a 64-molecule water system calculated with a 6-layer GNN that has a 6 angstrom cutoff distance would require the computation of 20,834 ghost atoms when using spatial partitioning (Musaelian_2023). Furthermore, unlike classical FFs, most MLIPs do not have a mature interface with LAMMPS, making spatial partitioning practically infeasible for the majority of MLIPs that have been developed.
@@
DeepMD is a short-range MLIP that has been applied to the simulation of 100 million atoms of water by 27,360 NVIDIA V100 GPUs on the Summit supercomputer, utilizing the spatial partitioning features within LAMMPS (Jia_2020_DP). The atomic system size was further extended to 10 billion atoms after further optimization of model tabulation, kernel fusion, and redundancy removal of the Deep Potential architecture (Guo_2022_DP).
@@
Allegro has been developed as a strictly local, E(3)-equivariant interatomic potential that features efficient parallelization through spatial partitioning due to its short-range design (Musaelian_2023). Because of this strict locality, Allegro demonstrated good scaling on large atomic systems – allegro_scaling used Allegro and LAMMPS to simulate a bulk Ag model with 100 million atoms, achieving 0.003 microseconds/atom-timestep using 128 NVIDIA-A100-80GBs.
@@
However, strictly local models experience key limitations. The need for efficient parallelization restricts their interaction range to only a few angstroms, preventing their use on systems that require the modeling of long-range interactions (Zhou2023; Song2024UNEP; Gong2025Bamboo; Cheng2025LES; anstine2023longrange). Furthermore, the short-range design prevents the MLIP’s application from simultaneously learning diverse chemical environments. As the short-range MLIP’s cutoff is often determined by the radial distribution function of one targeted material system, it is infeasible to determine a universal cutoff that efficiently works for many materials, which is becoming a common scenario with the increased interest in FPs. These problems raise the need for a simple, unified, and versatile API to parallelize MLIPs.
@@
In addition, graph-parallelization has been previously explored for training large GNN models anuroop. In comparison, the parallelization in MLIP inference poses a fundamentally different challenge compared to training. During training, the samples are almost always restricted to very small-scale chemical systems due to the computational complexity of acquiring labels. In training, the goal of graph-parallelization is to increase MLIP model sizes and batch sizes. During inference, graph parallelization is applied for the simulation of a single large chemical system. As a result, the application of graph-parallelization in large-scale MLIP simulations remains an open challenge.
@@ 3 Methods
Figure 1(a) denotes an overview of DistMLIP. Public MLIP models can be easily adapted to perform distributed large simulations with DistMLIP. The core infrastructure of DistMLIP is in the construction of graphs, subgraphs, and communication-related metadata.
@@ 3.1 Graph-Parallel Message Passing
After the material system is converted into a graph using a neighbor list construction algorithm, the graph can be partitioned into subgraphs for each device, as illustrated in Fig. 1(b). At each graph convolution, each node’s features are updated according to the edge and node features of its incoming neighbors. We can partition the nodes of the graph into disjoint sets, constructing graphs , where is the number of partitions. Each of the graph partitions are distributed to its own GPU. To accurately calculate the features after one graph convolution, we expand into , where consists of all nodes such that there exists an edge with . Formally,
@@
After each graph convolution, we transfer the border node and border edge features to and from each partition, as shown in Fig. 1(d). After this transfer process, each GPU has the most updated node and edge feature to begin the next convolution. This implementation is completely model-agnostic and can be applied to both conservative and direct force prediction MLIPs.
@@ 3.2 Distributing Atom Graphs
3.3 Distributing Higher-Order Graphs
Higher-order graphs, sometimes referred to as line graphs or bond graphs (for the three-body case), are frequently used in MLIPs to featurize higher-order interactions Choudhary_alignn_2021; deng2023chgnet; Zhang_2025_DPA3. Distributing the bond graph involves selecting all 1-hop and 2-hop neighbors of the pure atom graph nodes assigned to a partition. We then create an edge table mapping from node ids to edges originating from the node id pointing to a different node. By recursively traversing the table, we are able to create the bond graph for each partition in parallel. Border nodes within the bond graph are associated with the 1-hop edge neighbors of border edges within the atom graph – hence necessitating the inclusion of 2-hop neighbors. The parallel bond graphs thus contain the 1-hop neighbors of each pure bond graph node assigned to the partition. The complete procedure is found in algorithm 2 of the appendix.
@@ 3.4 Currently Supported MLIPs
Currently, we have implemented four widely used models in DistMLIP: 1) CHGNet (deng2023chgnet), an invariant MLIP that features both an atom graph as well as a bond graph, 2) TensorNet (tensornet; Ko2025Matgl), an atom graph-only, computationally-efficient, equivariant MLIP with performance on par with Nequip, 3) MACE (Batatia_2023), an atom graph-only equivariant MLIP that directly models many body interactions between atoms, and 4) eSEN (esen), an atom graph-only, smooth and equivariant MLIP. Specific architecture details and usage are found in appendix C and appendix I.